OpenStudy (anonymous):

If 8 = x and x = y, then 8 = y is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (unklerhaukus):

Reflexive A = A Symmetric A = B , B = A Commutative A*B = B*A Transitive A = B , B = C, A = C